### v1

# Still More Shades of Null: A Benchmark for Responsible Missing Value Imputation

Abstract.
We present Shades-of-Null, a benchmark for responsible missing value imputation. Our benchmark includes state-of-the-art imputation techniques, and embeds them into the machine learning development lifecycle. We model realistic missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR), to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test). Another key novelty of our work is that we evaluate imputers holistically, based on the predictive performance, fairness and stability of the models that are trained and tested on the data they produce.
We use Shades-of-Null to conduct a large-scale empirical study involving 20,952 experimental pipelines, and find that, while there is no single best-performing imputation approach for all missingness types, interesting performance patterns do emerge when comparing imputer performance in simpler vs. more complex missingness scenarios. Further, while predictive performance, fairness and stability can be seen as orthogonal, we identify trade-offs among them that arise due to the combination of missingness scenario, the choice of an imputer, and the architecture of the model trained on the data post-imputation. We make Shades-of-Null publicly available, and hope to enable researchers to comprehensively and rigorously evaluate new missing value imputation methods on a wide range of evaluation metrics, in plausible and socially meaningful missingness scenarios.
1. Introduction
As AI becomes more widely deployed into society, data — most importantly, openly accessible high quality AI-ready data — becomes a precious shared commodity. Among the factors affecting data quality is data missingness, a prevailing practical challenge of sustained interest to the data management, statistics and data science communities, and to the scientific community writ large (Little and Rubin, 1987; Rubin, 1976; Pigott, 2010; Mitra et al., 2023; Zhou et al., 2024; Shadbahr et al., 2023; Graham, 2009; Le Morvan et al., 2021; Zhang and Long, 2021a; Howell, 2007; Ilyas and Chu, 2019).
Specifically in data management, debates have been raging since the very inception of the field regarding whether — and how — to handle missing values, see  Date (1984) for a classic example. One difficulty, at the operational level, is that, while missing values are typically denoted with the token null, that may not always be the case. There could be hidden missing values in the data, like when ‘AL’ is selected as the state on a job application because it’s the first item in the pull-down list. Another difficulty, at the semantic level, is that a null can take on many meanings: it may signify that the value is unknown, inapplicable, intentionally withheld, etc.
In this paper, we do not concern ourselves with this semantic debate, and we also do not consider hidden missing values. Rather, we focus our attention on a very specific case. We are given a dataset $X$ (a single relation), in which the values of some features are missing, denoted by the marker null. The meaning of null is that the feature does take on a value in the real world (i.e., it is applicable) but that value is unobserved in $X$. Further, we intend to use $X$ in a machine learning (ML) setting: to train a model or to make predictions about the tuples in $X$ using a model that has already been trained and deployed. This model cannot deal with null values directly, and so these values must be imputed (filled in) as part of data pre-processing.
As our starting point, we will use Rubin’s missingness framework (Rubin, 1976) that, nearly 50 years since it was proposed, still remains the most popular approach to modeling missing data. Consider a dataset $X$ of $n$ samples, each with $p$ features, and an indicator $R$ such that $R_{i,j}=1$ when the value of the $j$’s feature of $X_{i}$ is missing: $X_{i,j}\text{is null}$, and $R_{i,j}=0$ when that the value is observed: $X_{i,j}\text{is not null}$. Rubin identified three data missingness scenarios:
Missing Completely at Random (MCAR)
Consider a dataset of job applicants that includes their most recent salary and the number of years of relevant work experience. The MCAR assumption holds, for example, when a job applicant’s most recent salary is missing from their record for purely administrative reasons that are uncorrelated with either their (potentially missing) salary or with their (observed) experience. More formally, the probability that a feature value is missing is independent of $X$: $P(R|X)=P(R)$.
Missing at Random (MAR)
Suppose that job applicants with fewer years of relevant work experience are more likely to withhold their most recent salary (in hopes of being offered a higher salary by the prospective employer), and that this can be explained by observed covariates (i.e., by the number of years of work experience). The MAR assumption holds in this example, where the probability that some feature value is missing depends only on the observed covariates $X_{\text{obs}}$, and not on the missing feature values: $P(R|X)=P(R|X_{\text{obs}})$.
Missing Not at Random (MNAR)
Suppose that a job applicant’s most recent salary does not correlate with their number of years of experience but rather is based on some combination of their geographic location and the results of a skills test, neither of which are captured in the data. Further, suppose that individuals whose most recent salary was low may be likely to withhold their salary information (again, in hopes of being offered a higher salary by the prospective employer). The MNAR assumption holds in this example, because missingness is correlated with the missing value itself, and it cannot be explained by the observed covariates (i.e., by the number of years of experience): $P(R|X)\not=P(R|X_{\text{obs}})$.
Missing value imputation (MVI)
Rubin’s framework has formed the basis of a large body of technical work on missing value imputation (MVI), which has been covered in several excellent and comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019). Broadly speaking, MVI approaches fall into several categories: (1) statistical, such as imputing missing observations with the median or mode (Schafer and Graham, 2002); (2) learning-based impute-then-classify, which use the observed data to iteratively impute and learn missing values, using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); and (3) joint data cleaning and model training (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s seminal multiple imputation framework (Rubin, 1987).
Beyond Rubin’s framework: Mixing scenarios and dealing with missingness shift.
Rubin’s framework, while clean and amenable to analysis, does not fully capture real-world missingness. First, it has been argued that the assumptions behind MCAR rarely hold in practice, and that most real-world data exists on a continuum between MAR and MNAR, depending on how that data was collected (Graham, 2009). Second, while it is useful to distinguish between missingness scenarios based on their statistical properties to draw valid inferences, we are likely to encounter multi-mechanism missingness in practice, where several scenarios may co-exist in the same dataset, affecting subsets of its features, tuples, or both (Zhou et al., 2024). For example,  Mitra et al. (2023) introduce the notion of a data missingness life cycle and posit that data pre-processing steps such as assimilating data from different sources or from different heterogeneous populations (each with their own missingness patterns) can result in ‘structured missingness’ that cannot be represented using Rubin’s framework. Third, it has been observed that, in the context of data-centric AI, the assumptions that hold about missingness in the training data may no longer hold post-deployment, a phenomenon termed missingness shift (by analogy with data distribution shift) (Zhou et al., 2023).
Evaluating MVI techniques.
Despite an abundance of MVI techniques using different model architectures being proposed each year, making meaningful progress on learning from incomplete data ultimately comes down to how realistically we model missingness: our proposed methods are only as good as the evaluation scenarios we rigorously test them on. If our evaluation scenarios do not realistically capture the kinds of missingness that we encounter in practice, then our data cleaning methods are unlikely to be effective. Further, newly proposed MVI techniques are often only evaluated on the correctness of the imputation (imputation quality) and the correctness of the predictive model (accuracy or F1). However, in recent years, the scientific community is starting to uncover trade-offs between MVI efficacy and other normatively important dimensions of model performance, including algorithmic fairness (Guha et al., 2024; Feng, 2023), and stability and robustness (Zhou et al., 2024), beyond accuracy. We elaborate on this next.
Missingness as a form of pre-existing bias.
Consider again the job applicant screening example, and suppose that the applicants’ gender and age are among the features. Female job applicants who suspect that they have been receiving lower salaries than their male colleagues, despite having comparable qualifications and job experience, may withhold salary information more frequently then men, hoping to narrow the gender wage gap in their next position. As a result there will be more missing salary values for women, and the probability that the value is missing depends on the observed covariate, gender. This situation is consistent with MAR missingness, and it points to pre-existing bias — a situation where data reflects a history of societal discrimination (Friedman and Nissenbaum, 1996). For our final example of pre-existing bias, suppose that disability status is among the features. Job applicants who live with a disability are more likely to omit a value for this feature. Assuming that this feature does not correlate with the other features, this situation is consistent with MNAR, with missingness itself serving as a proxy for disadvantage.
Summary of Contributions
We implemented an experimental benchmark called Shades-of-Null to rigorously and comprehensively evaluate state-of-the-art MVI techniques on a variety of realistic missingness scenarios (including single- and multiple-mechanism missingness and missingness shift), on a suite of evaluation metrics (including fairness and stability), in the context of data pre-processing in a machine learning pipeline.
Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 8 MVI techniques on 6 benchmark datasets using 5 model types, running a total of 20,952 pipelines, and is the first study of such scale in the missing data domain to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, admissions, and healthcare, where missingness is socially salient. Mitigating social harm such as algorithmic discrimination is a leading concern in these domains, and we evaluate the impact of MVI approaches on downstream model fairness and stability (which have been understudied in the context of missing data), in addition to classically studied imputation quality and model correctness metrics.
As an additional engineering contribution, while developing the Shades-of-Null benchmark, we improved several state-of-the-art MVI techniques, see Appendix A.2 for details.
We make Shades-of-Null publicly available [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively and rigorously evaluate new missing value imputation methods on a wide range of evaluation metrics, in plausible and socially meaningful missingness scenarios.
2. Related Work
Missing value imputation (MVI) techniques
Learning-based approaches have become increasingly popular, and include k-Nearest Neighbors, Decision Trees, Support Vector Machines, clustering, and ensembles (Batista and Monard, 2003; Lin and Tsai, 2020; Hasan et al., 2021).  Zhou et al. (2024) and Liu et al. (2023) review Deep Learning-based approaches (Variational Auto-Encoders and GANs) and representation learning (Graph Neural Networks and diffusion-based methods). Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and Expectation Maximization (Pigott, 2010) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022). According to (Feng, 2023; Le Morvan et al., 2021), impute-then-classify is suboptimal compared to multiple imputation because missingness itself can be informative and is lost before model training.
Beyond impute-then-classify, the data management community has proposed holistic methods like CPClean (Karlaš et al., 2020) and ActiveClean (Krishnan et al., 2016), that jointly perform data cleaning and model training, deriving from the Multiple Imputation framework (Rubin, 1987). These methods detect and repair a variety of errors including outliers, mislabels, duplicates, and missing values, and hence are less directly optimized to model missingness, instead focusing on improving data quality holistically. BoostClean (Krishnan et al., 2017) aims to reduce the human effort in error repair by learning efficiently from a few gold standard annotations (from a human oracle).
Evaluating MVI techniques
Empirical studies have been overwhelmingly focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008). Further, most proposed methods only perform direct evaluation, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022). Very few studies use both direct and indirect evaluation, and even studies that include indirect evaluation only evaluate predictive accuracy (Lin and Tsai, 2020). This leaves a significant research gap, specially for desiderata such as stability and self-consistency, which have recently been shown to be highly sensitive to data-centric interventions (Black and Fredrikson, 2021; Black et al., 2022; Cooper et al., 2023), and unfairness, which can be worsened by improperly chosen data cleaning steps (Guha et al., 2024). Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
Notably, and worryingly, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics (including size and correlation between variables) (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
Fairness and missingness
There has been some recent interest in studying the social harm that can come from poorly chosen missing value imputation techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024). Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022) have worked with the COMPAS and Adult datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021). These datasets do not contain naturally-occurring missing values, and experimental studies employ randomly-generated missingness: usually by sampling or using a fixed set of columns, and randomly picking rows in which to replace values with null. We critique this approach, since detecting and mitigating unfairness requires broader socio-technical thinking, such as having higher rates of missingness for minority groups and in features that are highly correlated with sensitive attributes (called proxy variables in the fairness literature) (Caton et al., 2022).
A notable exception is  Martínez-Plumed et al. (2019), who map social mechanisms such as prejudicial access and self-reporting bias to different categories of missingness such as missing-by-design and item non-response. Further, they analyse the correlations between all features to study the effect of different kinds of missingness. We use a similar methdology to simulate realistic missingnesses in this work. Importantly, however, we find conceptual limitations in their framing of algorithmic fairness. Martínez-Plumed et al. (2019) write:“The surprising result was to find that, […] the examples with missing values seem to be fairer than the rest.” Asserting that some rows of data are more or less fair does not make sense, in that samples (e.g., job applicants) do not demonstrate the property of being fair. It is the model (e.g., used to make hiring decisions) that asserts the property of being fair or not, by the inclusion or exclusion of certain samples amongst those given a positive outcome. We re-interpret the findings of  Martínez-Plumed et al. (2019) to say that excluding samples with missingness can increase model unfairness, which is further evidence against the use of deletion as a strategy for dealing with missing data.
Zhang and Long (2021b) evaluate MVI methods on imputation fairness, measured as the imputation accuracy difference. They find that (1) higher missingness disparity leads to higher imputation unfairness; (2) higher rates of missingness lead to higher imputation unfairness; (3) more imbalance in the data (or larger disparity in the sample sizes across groups) leads to higher imputation unfairness, since there is an imbalance of observed information; (4) the same imputation method has different prediction fairness under different missingness mechanisms. Their evaluation is limited to randomly generated null values in COMPAS. We extend their analysis of fairness trade-offs to additional datasets and missingness scenarios.
Zhang and Long (2022) introduce the notion of imputation fairness risk as the disparity in imputation quality for different demographic groups, and provide bounds on the fairness risk of “correctly specified” imputation methods. While there is very little theoretical work in this space, and this is a commendable contribution, we call into question the broader implications of this result: it is well-known that imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a). Put differently, a classifier built on poor imputation quality can still give good downstream performance (Shadbahr et al., 2023). With this in mind, we ask the question: Does minimizing imputation unfairness minimize model unfairness? Our empirical findings indicate that they do not, as is discussed in Sections 5.4.
Missingness and stability
We are not aware of any work investigating the effect of missing value imputation on model stability.
3. Benchmark Overview
3.1. Methodology for Simulating Missingness
In this work, we start with datasets in which there are no missing values, and then simulate missingness. We make this choice to be able to evaluate the effectiveness of missing value imputation scenarios with respect to the ground truth. Our methodology for simulating missingness is based on evaluation scenarios, defined by the missingness mechanism during train and test, shown in Table 1: (1) single-mechanism missingness, injected similarly into train and test (S1 - S3); (2) single-mechanism missingness, injected differently into train and test (missingness shift) (S4 - S9); and (3) missingness is mixed, to include all three missingness mechanisms, and is injected similarly into train and test (S10).
Table 1. Table 1: Evaluation Scenarios
Train     Test
Scenario MCAR MAR MNAR MCAR MAR MNAR
S1
S2
S3
S4
S5
S6
S7
S8
S9
S10
For each dataset in our benchmark (see Section 3.5), we devised dataset-specific missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR). These scenarios were crafted to be realistic, and were guided by the correlation coefficients between all covariates and the target variable, as well as by feature importance derived using scikit-learn’s built-in functionality, following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019). Furthermore, we selected columns that may plausibly contain missingness. For example, blood pressure or cholestrol in diabetes is expected to always be observed, but family history or physical activity might be omitted or withheld. Table 3 presents missingness scenarios for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Details for other datasets are deferred to the Appendix B.2.
3.2. Missing Value Imputation (MVI) Techniques
As discussed in Section 2, there is a large number of competitive MVI techniques in the literature. For our study, we picked 8 techniques, from 5 broad categories in the taxonomy of Emmanuel et al. (Emmanuel et al., 2021), namely: (1) baseline: deletion; (2) statistical techniques: median-mode and median-dummy; (3) machine learning-based techniques: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) deep learning-based techniques: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); and (5) multiple imputation: boostclean (Krishnan et al., 2017). See Appendix A.1 for details. We limit our analysis to one joint-cleaning-and-training approach, boostclean (Krishnan et al., 2017), and leave the exploration of other methods, such as ActiveClean (Krishnan et al., 2016) and CPClean (Karlaš et al., 2020) to future work.
3.3. Evaluation Metrics
Following (Lin and Tsai, 2020; Hasan et al., 2021), we evaluate performance of MVI techniques in two ways: directly using imputation quality metrics and indirectly based on downstream model performance metrics.
3.3.1. Imputation Quality Metrics
Shadbahr et al. (Shadbahr et al., 2023) report that classically used imputation quality metrics are uncorrelated to downstream model performance, and that distributional metrics perform better. To confirm or refute this claim, we use a mix of discrepancy and distributional metrics.
To assess agreement with true values, we compute Root Mean Square Error (RMSE) for numerical and F1 score for categorical attributes. To assess distributional alignment, we compute KL-divergence for both numerical and categorical attributes, for the imputed values as well as for the entire training and test sets. We compute KL-divergence as the Shannon entropy of the probability distributions of the true and predicted values. For categorical attributes, we obtain the probability distributions using the value_counts method with normalization from pandas. For numerical attributes, we first estimate probability density functions (PDFs) using Gaussian kernel density estimation from the scipy package, and then evaluate these estimated PDFs on a set of 1,000 random points to derive the probability distributions. Finally, to assess imputation fairness (Zhang and Long, 2021b, 2022), we compute F1 score difference, RMSE difference, and KL divergence difference between privileged and disadvantaged groups.
3.3.2. Model Performance Metrics
To evaluate model performance we utilize Virny (Herasymuk et al., 2024), a comprehensive software library that computes fairness, stability, and accuracy metrics for a set of initialized models, on several single-attribute and intersectional subgroups of interest. Specifically, to assess the impact of MVI techniques on model correctness, we report the F1 score because it is a more reliable metric than accuracy for imbalanced data. For evaluating model stability, we report average Label Stability (Darling and Stracuzzi, 2018; Khan et al., 2023) over the full test set (closely related to the self-consistency metric from Cooper et al. (Cooper et al., 2023)), computed per-sample for binary classification Label Stability = $\frac{|B_{+}-B_{-}|}{B}$, where $B_{+}$ is the number of times the sample is classified into the positive class and $B_{-}$ is the number of times the sample is classified into the negative class, and $B=B_{+}+B_{-}$ models are trained by bootstrapping over the train set. We set $B=50$ in all of our experiments.
Lastly, we compute standard fairness metrics based on group-specific error rates, namely True Positive Rate Difference (TPRD), True Negative Rate Difference (TNRD), Selection Rate Difference (SRD), and Disparate Impact (DI). (Note that DI computes the ratio of selection rates, but we refer to it as DI as is standard in the literature (Feldman et al., 2015).)
$TPRD=\frac{TP_{dis}}{TP_{dis}+FN_{dis}}-\frac{TP_{priv}}{TP_{priv}+FN_{priv}}$
$TNRD=\frac{TN_{dis}}{TN_{dis}+FP_{dis}}-\frac{TN_{priv}}{TN_{priv}+FP_{priv}}$
$SRD=\frac{TP_{dis}+FP_{dis}}{N_{dis}}-\frac{TP_{priv}+FP_{priv}}{N_{priv}}$
$DI=\frac{TP_{dis}+FP_{dis}}{N_{dis}}\div\frac{TP_{priv}+FP_{priv}}{N_{priv}}$
3.4. Shades-of-Null Benchmark Architecture
Figure 1 summarizes the architecture of the Shades-of-Null benchmark. Our goal when developing this benchmark was to support a large-scale evaluation of the impact of different missingness scenarios (Section 3.1) and missing value imputation techniques (Section 3.2) on different aspects of performance of machine learning models (Section 3.3). This benchmark is novel in several ways. First, it handles both single-mechanism and mixed missingness scenarios, including missingness shift. Second, it supports flexible evaluation of imputation quality and model performance — including accuracy, fairness and stability — for multiple sensitive attributes and protected groups, as well as their intersections.
The central component of the architecture in Figure 1 is the benchmark controller that executes the specified evaluation scenarios and applies error injectors to input datasets, accordingly. It then imputes missing values and preprocesses the imputed datasets using standard scaling for numerical features and one-hot encoding for categorical features. The controller then conducts ML model hyperparameter tuning. Finally, the evaluation module assesses MVI and ML model performance, both overall and for specific subgroups. To facilitate comprehensive model profiling, the evaluation module utilizes the Virny software library (Herasymuk et al., 2024), measuring correctness, fairness and stability. We store experimental results in MongoDB.
Figure 1. Shades-of-Null benchmark architecture
Table 2. Table 2: Dataset Information
Name Domain Size Sensitive Attributes
diabetes healthcare 905 sex
german finance 1,000 sex, age
folk finance 15,000 sex, race
law-school education 20,798 sex, race
bank marketing 40,004 age
heart healthcare 70,000 sex
Shades-of-Null incorporates two optimizations to streamline the execution of multiple experiments. Firstly, it separates missing value imputation and model evaluation stages. This enables the preservation of imputation results, including imputed datasets stored in a file system and imputation performance metrics stored in a database, facilitating their reuse during the subsequent model evaluation stage. Consequently, redundant application of the same MVI on identical datasets with identical seeds for the pipeline is avoided when assessing different models. Secondly, the experimental pipeline within the benchmark is tailored to evaluate MVI and ML models simultaneously on multiple test sets, with varying error rates and missingness types. As both MVI and ML models only perform inference on test sets, the running time for executing a pipeline with one training set and one test set is close to that of a pipeline with one training set and multiple test sets.
3.5. Datasets and Tasks
A summary of the datasets used in this study is given in Table 2.
Table 3. Table 3: Missingness scenarios for an error rate of 30% for diabetes. SoundSleep is a numerical column; Family__\__Diabetes, PhysicallyActive and RegularMedicine are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR SoundSleep, Family__\__Diabetes, PhysicallyActive, RegularMedicine N/A 0.3 0.3
MAR Family__\__Diabetes, RegularMedicine Sex 0.2 (female) 0.1 (male)
PhysicallyActive, SoundSleep Age 0.2 ($\geq 40$) 0.1 ($<40$)
MNAR Family__\__Diabetes Family__\__Diabetes 0.25 (yes) 0.05 (no)
RegularMedicine RegularMedicine 0.2 (yes) 0.1 (no)
PhysicallyActive PhysicallyActive 0.25 (none, $<\frac{1}{2}$ hour) 0.05 ($>\frac{1}{2}$ hour, $>1$ hour)
SoundSleep SoundSleep 0.2 ($<5$) 0.1 ($\geq 5$)
(a) Correlation with label
(b) Feature importance
Figure 2. EDA for designing missingness scenarios in diabetes.
diabetes [^2] https://www.kaggle.com/datasets/tigganeha4/diabetes-dataset-2019 (Tigga and Garg, 2020) was collected in India through a questionnaire including 18 questions related to health, lifestyle, and family background. A total of 952 participants are characterized by 17 attributes (13 categorical, 5 numerical) and a binary target variable that represents whether a person is diabetic. Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
german [^3] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data (Hofmann, 1994) contains records of creditworthiness assessments, classifying individuals as high or low credit risks. It contains information on 1,000 individuals characterized by 20 attributes (13 categorical, 7 numerical). Here, sex and age are the sensitive attributes, with “female” and ¡25 as the disadvantaged groups.
folk (Ding et al., 2021) was derived from US Census surveys conducted in 2014-2018, and is associated with several tasks. We selected ACS Income, a binary classification task to predict whether a person’s annual income is above \$50,000. We use 2018 data from Georgia, with 15,000 tuples and 10 attributes (8 categorical, 2 numerical). Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
law-school [^4] https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage  (Wightman, 1998) was gathered through a survey conducted by the Law School Admission Council (LSAC) across 163 law schools in the US in 1991, and contains law school admissions records of 20,798 applicants, characterized by 12 attributes (6 categorical, 6 numerical). The task is to predict whether a candidate would pass the bar exam. Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
bank [^5] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013. It contains information on 40,004 potential customers, with 16 attributes (10 categorical, 6 numerical) and a binary target that represents whether the individual signed up for the product. Here, age is the sensitive attribute, with ¡25 and ¿60 as the disadvantaged group.
heart [^6] https://www.kaggle.com/datasets/sulianova/cardiovascular-disease-dataset consists of patient measurements with respect to cardiovascular diseases, wiht information of 70,000 individuals characterized by 11 attributes (6 categorical, 5 numerical), and a target variable that denotes the presence of a heart disease. Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
4. Single-Mechanism Missingness
In our first set of experiments, we compare the performance of MVI techniques in single-missingness scenarios (S1-S3 in Table 1), with the same fraction of errors (30%) in train and test.
For correctness, we report results for F1, and defer additional results to Appendix C.2. For fairness, we report metrics based on an intersectional group (e.g., sex&race) for datasets with two sensitive attributes, and a single group (e.g., age) for datasets with one sensitive attribute. For stability, we used a bootstrap of 50 estimators, each seeing a random 80% of the train set (Efron and Tibshirani, 1994). Higher values of F1 and label stability are better, and values of TPRD close to zero are better.
We evaluate 5 ML types: Decision Tree (DT), Logistic Regression (LR), Gradient Boosted Trees (LGBM), Random Forest (RF), and a Neural Network (historically called the Multi-Layer Perceptron or MLP). Different models may perform best on different datasets. In Figures 34 and 5, we report on the best-performing models (based on their F1 score) for each dataset, and make it explicit that we are comparing across different model types. Additionally, we compare performance against a model that is trained on the clean train set using the same model architecture.
4.1. Correctness of the Predictive Model
As shown in Figure 3, the efficacy of MVI techniques, and the downstream effect on the predictive performance of the model, according to F1, depends significantly on the characteristics of the dataset and the missingness mechanism.
All the methods we evaluated are competitive for all missingness mechanisms on heart and law-school. boostclean, which uses multiple imputation (MI), is competitive only on small datasets (diabetes and german), and only under MCAR. It performs particularly poorly on folk, with as much as a 0.8 drop in F1 compared to other methods. We discuss this unexpected performance of MI further in Section 7. In line with conventional wisdom, deletion significantly worsens predictive performance for small datasets. For diabetes (the smallest dataset), F1 drops by 0.06 for MCAR, 0.08 for MAR and 0.1 for MNAR, compared to ML model trained on clean data. Under MCAR, miss-forest (ML-based) and auto-ml (DL-based) are the best-performing techniques. Under MAR and MNAR there is no clear best-performing technique. For example, under MAR datawig ( DL-based) shows good performance, but simple median-mode and median-dummy imputers are also competitive. This underscores the need to evaluate novel DL-based and ML-based approaches more holistically (on a variety of missingness mechanisms) to ensure that they justify the additional training overhead and complexity they introduce, compared to simple statistical methods.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 3. F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are ordered in increasing order by size. The blue line indicates the performance of a model trained on clean data.
4.2. Fairness of the Predictive Model
Next, we evaluate the fairness of the downstream ML model for different MVI techniques. We present TPRD in Figure 4, with results for other fairness metrics in Appendix C.2. The effect of MVI on fairness (error-disparity) is highly dataset specific, and it is strongly correlated with fairness of the model trained on clean data (no missingness), corroborating the findings of Guha et al. (Guha et al., 2024).
On heart and german, we observe negligible disparity for the model trained on clean data, and MVI techniques do not further exacerbate it. On folk and bank, we see large disparities (close to 20% and 10%, respectively, for model trained on clean data). Interestingly all MVI techniques preserve or improve fairness (reduce disparity) in all missingness settings, with boostclean showing lowest disparity. On diabetes and law-school (which, for the model trained on clean data, show disparities close to 10% and 8%, respectively), MVI techniques such as boostclean, clustering and deletion can worsen fairness, with the largest effect under MNAR.
Wang and Singh (2021) posit that models will exhibit more unfairness under MAR and MNAR compared to MCAR, but we only find weak empirical evidence towards this, even for techniques like deletion. A nuance here is in the way we set up the missingness scenarios, described in detail in Section 3.1. We took a broad socio-technical lens and designed non-response scenarios that are realistic — including MAR scenarios where people from disadvantaged groups withhold information that could hurt their chances of getting the desired outcome. Hence, dropping these rows might in fact improve “fairness” by removing feature values that contribute negatively to outcomes for the disadvantaged group. With this intuition, deletion can in fact improve fairness under MAR and MNAR, as observed on bank and diabetes.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 4. True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are desirable. Datasets are ordered in increasing order by size. Blue line shows performance of the model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 5. Label Stability of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are desirable. Datasets are ordered in increasing order by size. The blue line indicates performance of the model trained on clean data.
4.3. Stability of the Predictive Model
Next, we compare the stability of the downstream model trained using different MVI techniques. Results in Figure 5 show that, as expected, stability depends most strongly on dataset characteristics (mainly size), then on imputation strategy.
All imputers preserve or marginally improve stability on bank and heart (largerst datasets, with 40k and 70k tuples, respectively), for all missingness mechanisms. Even deletion does not worsen stability in this case since there is sufficient data even after dropping rows with nulls, and clustering marginally improves stability over the model trained on clean data, for all missingness types.
On folk and law-school (medium-sized datasets, with 15k and 20k tuples, respectively), most MVI techniques are competitive, except for boostclean, which shows a 5% drop in label stability compared to the model trained on clean data. boostclean does joint cleaning and model training, and is expected to be sensitive to small changes in the training data, as validated by our findings.
On german and diabetes (about 1K samples each) deletion, clustering and boostclean are the worst performing methods on all missingness mechanisms, with label stability 8-10% lower compared to the model trained on clean data. miss-forest and auto-ml are the best performing methods, and match the stability of the model trained on clean data for all missingness types.
4.4. Imputation Quality and Fairness
To better understand the impact of different MVI techniques on model performance, we measure the imputation quality of each imputer before training ML models on the imputed training set. The quality of imputation is evaluated across three dimensions: correctness (using F1 for categorical features and RMSE for numerical features), distributional alignment (using KL divergence of imputed values), and fairness (measuring differences in F1, RMSE, and KL divergence between disadvantaged and privileged groups). The higher values of F1 are the better, the lower values of RMSE and KL divergence are the better, and values of imputation fairness closer to zero are the better.
To streamline exposition, we selected one categorical and one numerical feature for each dataset based on their Spearman correlation with the target. However, our analysis takes into account the imputation quality measured on all features with nulls. Importantly, we do not assess imputation quality for boostclean, since this is a joint cleaning and model training method that does not provide imputed values, which we could use to measure imputation quality.
4.4.1. Correctness of imputation
Figures 18 and 19 in Appendix C.1 demonstrate the correctness of imputation of all MVI techniques, except boostclean, for all datasets. From Figure 18, we can notice that one of the ML and DL-based imputers (miss-forest, auto-ml, datawig) is the best in terms of F1 for categorical columns under MCAR and MAR missingness mechanisms. Additionally, Figure 19 shows that all the three ML and DL-based imputers are dominant in terms of RMSE for numerical columns compared to statistical MVI techniques. However, these significant differences in F1 and RMSE do not translate into substantial differences in downstream model performance, as measured by F1 in Section 4.1. This observation further supports our statement made in Section 5.4 that “imputation quality metrics are not very meaningful because a bad data imputation method can be overcome by a strong learner, and so bad performance on imputation quality does not necessarily mean than the MVI will lead to poor predictive performance.”
4.4.2. Distributional alignment of imputation
Figures 20 and 22 in Appendix C.1 illustrate the distributional alignment of imputation of all MVI techniques except boostclean, for all datasets. We find that ML and DL-based imputers, except clustering, outperform statistical approaches in terms of distributional alignment across all missingness mechanisms. The plots clearly demonstrate that, as we highlight in Section 5.4, distributional metrics often serve as more reliable indicators than discrepancy metrics. This is because the distributional alignment of imputed features can significantly influence model training outcomes, emphasizing that assessing only the correctness of imputation is often insufficient.
4.4.3. Fairness of imputation
Figures 21, 23, 24, and 25 in Appendix C.1 show the fairness of imputation of all MVI techniques, except boostclean, for all 6 datasets. From Figures 21 and 23, we observe that statistical MVI techniques are prone to worsen imputation fairness rather than improve it under all missingness mechanisms. Further, comparing imputation fairness with the fairness of the downstream predictive model, discussed in Section 4.2, we can state that good imputation fairness may not lead to model fairness downstream. For example, for the folk dataset, miss-forest was better than median-mode in terms of imputation fairness (F1 score difference of 0.08), yet the downstream model trained on miss-forest-imputed data was less fair than the model trained on median-mode-imputed data (0.03 TRPD).
5. Multi-mechanism Missingness
We now evaluate different MVI techniques under multi-mechanism missingness (Zhou et al., 2024), when MCAR, MAR and MNAR co-exist in both train and test data (S10 in Table 1). This is a more realistic setting that is likely to be encountered in practice (Graham, 2009). We inject 10% of nulls from each mechanism into both train and test sets, for a total of 30% missingness, as in S1-3. In Section 4, we reported results for the best performing ML models for each dataset, which was the Random Forest for diabetes. In this experiment, we focus on diabetes and dive deeper into the performance of different model types.
(a) F1
(b) True Positive Rate Difference (Fairness)
(c) Label stability
Figure 6. Performance of MVI techniques under multi-mechanism missingness on the diabetes dataset, for different model types: Decision Tree (dt__\__clf), Logistic Regression (lr__\__clf), Gradient Boosting (lgbm__\__clf), Random Forest (rf__\__clf) and Neural networks (mlp__\__clf). Boostclean is a joint cleaning and classification method, reported as a separate model type.
(a) F1 (categorical), RMSE (numerical)
(b) KL divergence for all columns
Figure 7. Imputation quality on all columns with mixed missingness for diabetes
(a) F1 disparity (categorical), RMSE disparity (numerical)
(b) KL divergence disparity for all columns
Figure 8. Imputation fairness on all columns with mixed missingness for diabetes
5.1. Correctness of the Predictive Model
Recall from Section 4.1 that miss-forest was the best performing MVI method for diabetes, under all single-mechanism types, and it was the only method that performed comparably to performance of the model trained on clean data under MNAR. According to Figure 6 (a), miss-forest remains one of the best-performing MVM techniques in terms of F1, for all model types. Further, under single-mechanism missingness, auto-ml (DL-based) was only competitive under MCAR and MAR, however, under multi-mechanism missingness, its F1 is comparable to that of miss-forest, for all model types.
Notably, while boostclean was not the best performing under any single-mechanism missingness, it is highly competitive under multi-mechanism missingness, and even outperforms miss-forest and auto-ml for Decision Trees (DT), Logistic Regression (LR) and Gradient Boosting (LGBM), indicating the competitive edge of joint cleaning and training over impute-then-classify, under complex missingness scenarios (Feng, 2023; Le Morvan et al., 2021). Finally, as for single-mechanism missingness, deletion continues to show the worst performance, with the largest effect of nearly a 15% drop in F1 compared to the model trained on clean data for LGBM.
5.2. Fairness of the Predictive Model
Recall from Section 4.2 that all MVI methods worsened fairness for the RF model for diabetes under single-mechanism missingness, with clustering being the worst-performing for all missingness types. From Figure 6 (b) we see that the effect of MVI on fairness is highly ML model-specific: for example, clustering improves fairness for LR but worsens it for all other model types, including RF. Note that LR has the worst unfairness (highest TPRD) when the model is trained on clean data, even as compared to deletion with other ML model types. miss-forest, datawig, and auto-ml preserved fairness under single-mechanism missingness, and continue to show good performance under multi-mechanism missingness, for all model types.
5.3. Stability of the Predictive Model
Recall from Section 4.3 that deletion, clustering and boostclean worsened stability under all three single-mechanism missingness types. In Figure 6 (c), deletion and clustering continue to show poor performance for all model types. Interestingly, boostclean (multiple imputation with joint cleaning and training) is competitive under multi-mechanism missingness: it shows higher label stability than all other MVI approaches for the DT, and is competitive with the best performing imputers for MLP and LGBM.
Simple imputers like median-mode and median-dummy— which were competitive on single-mechanism missingness — are no longer competitive under multi-mechanism missingness, while ML- and DL-based approaches like miss-forest, auto-ml and datawig continue to show good stability.
5.4. Imputation Quality and Fairness
Since multi-mechanism missingness has never been empirically evaluated before (to the best of our knowledge), we additionally report performance on imputation quality metrics in Figure 7. We do not include boostclean since it does not provide imputed values, which we could use to assess imputation quality. In line with  Shadbahr et al. (2023), we find that distributional metrics are more indicative of downstream performance than discrepancy metrics.
deletion, median-mode, and clustering all perform poorly on both discrepancy and distributional metrics, and are also the worst performing in terms of downstream model correctness (for F1). datawig, however, has high error on F1 and RMSE, but not on KL, and this is indicative of its good performance of the downstream predictive model. Further, as part of this work we assert that correctness of downstream models should be evaluated alongside their fairness and stability. We find that imputation quality metrics — especially discrepancy metrics — are not correlated with fairness and stability downstream.
On imputation fairness, auto-ml and miss-forest show the lowest disparity, but they show residual predictive unfairness, as shown for Figure 6. Further, miss-forest has better imputation fairness than median-mode, but leads to worse unfairness of the downstream models. For example, as can be seen in Figure 8 (b), miss-forest has much better imputation fairness than median-mode (0.20 F1 score difference). Yet, as demonstrated in Figure 6 (b), median-mode leads to better fairness of the RF model compared to miss-forest (0.03 TRPD). These observations, once again, corroborate what we observed for for single-mechanism missingness (Section 4.4.3): assessing imputation fairness without considering the fairness of the downstream model is often insufficient.
6. Missingness Shift
Next, we evaluate the effect of missingness shift on the correctness, fairness, and stability of the downstream model. For that, we vary missingness mechanisms and missingness rates in train and test sets separately, first by holding the fraction of nulls in the test set constant (at 30%) and varying the fraction of nulls in the train set (10%, 30% and 50%). Then, we hold the fraction of nulls in the train set constant (at 30%) and vary the fraction of nulls in the test set ($10\%$, $20\%$, $30\%$, $40\%$ and $50\%$). Note that we have fewer settings for train set fractions because varying the test set is less computationally demanding (see Section 3.4).
We report results on diabetes. See Appendix D and GitHub repository for additional results on other datasets, with fixed and variable train and test missingness fractions.
6.1. Correctness of the Predictive Model
Figure 9. F1 of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Figure 10. F1 of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Train set missingness. As expected, according to Figure 9 the F1 of the predictive model depends both on the fraction of train nulls, and the missingness mechanism. Under MCAR, for low rates of missingness (10% and 30%), the F1 is constant, and then starts decreasing with more nulls (50%). Under MNAR, with deletion the decrease in F1 is monotonic, and this is in line with conventional wisdom. Under MNAR, the missingness depends on the value itself. This means that when values are missing, they cannot be inferred by other values in the observed data. For this reason, training under MNAR is very sensitive to the train missingness rate.
Interestingly, correctness of ML and DL methods (miss-forest, auto-ml, datawig), and of multiple imputation (boostclean) is relatively insensitive to train set missingness rate. Even at 50% error these methods have sufficient data to learn the missingness pattern reasonably well. This is a striking result given that diabetes was the smallest dataset in our study with only 1,000 samples. Another interesting observation is that clustering is the only method that improves on F1 as the fraction of nulls in the train set increases, for all types of missingness as well as under missingness shift. However, at high rates of training error, even simple imputers like median-mode and median-dummy outperform clustering.
Test set missingness. Figure 10 shows the F1 of the RF model on diabetes, as a function of test set error rate under different single-mechanism missingness (S1-3) and missingness shift (S4-9). We observe that most methods are relatively robust to change in test missingness rates, including also under missingness shift. An important exceptions are scenarios S5 and S7 (Table 1), when train data is missing under MCAR or MAR, while test data is missing under MNAR. For these scenarios, most MVI methods show a steady decrease in F1 with increasing test set missingness rate.
miss-forest is very robust, and shows comparable F1 for all missingness scenarios (including shifts) and test missingness rates. On the other hand, deletion is the most sensitive to the test error rate, and shows the poorest performance on MNAR, even without missingness shift. Our results validate conventional wisdom that MCAR and MAR are easier to model and impute than MNAR and that, in general, increasing the rate of missingness in the test set decreases predictive performance (Shadbahr et al., 2023).
6.2. Fairness of Predictive Model
Figure 11. True Positive Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Figure 12. True Positive Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Train set missingness. Figure 11 shows that all methods are sensitive to the train missingness rate when it comes to fairness. For example, miss-forest has stable performance under MNAR train. However, for MCAR and MAR training, the most fair setting is when the error rate in train and test is the same (30%). For larger (50%) and smaller (10%) train error rates, the model shows increased disparity (unfairness). Most ML and DL-based approaches such as auto-ml, datawig, and miss-forest exhibit this behavior, indicating that parity in train and test error rates is one of the most important factors for fairness.
Test set missingness. Figure 12 demonstrates that the trend depends most strongly on test missingness type: deletion under MAR improves fairness as the test error rate increases (for the same reason discussed in Section 4), whereas under MNAR it steadily worsens it. Under MCAR, fairness decreases when the test error rate exceeds the train error rate (30% in train and 50% in test). The miss-forest method is highly robust, with fairness worsening only marginally (by 2%) under MNAR test missingess. In contrast, ML and DL-based methods like clustering, datawig and auto-ml show worse fairness with an increase in the test error rate.
6.3. Stability of Predictive Model
Train set missingness. Figure 30 in Appendix D.1 illustrates that, under MCAR and MAR train missingness, most methods with the exception of boostclean, deletion, and clustering show good stability, and are relatively insensitive to the increase in the train error rate. deletion shows the largest drop in stability with an increase in the train error rate, and this is in line with conventional wisdom as dropping rows with nulls decreases the effective train set size, and less training data makes models more unstable.
Techniques like datawig and boostclean show worse stability under high train error rates, and this effect is the most pronounced under MNAR. miss-forest is, once again, highly robust, and maintains good stability (comparable to that of the model trained on clean data) under all missingness scenarios, rates and shifts.
Test set missingness. Figure 31 in Appendix D.1 shows that the test missingness rate has very little effect on the stability, for most MVI approaches, with the exception of clustering, which shows higher instability at higher rates of test error, with the largest effect in scenario S7 (MAR train, MNAR test set).
7. Summary of Experimental Findings
Deletion is the worst strategy. Building on evidence in the scientific community(Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our findings indicate that, in addition to hurting predictive performance, deletion also worsens stability and fairness, arguing against the use of this method in practice.
Multiple imputation shows mixed results.
There is conflicting evidence in the literature on the performance of multiple imputation (MI), and our empirical findings are also mixed and somewhat unexpected. Feng (Feng, 2023) and Le Morvan et al. (Le Morvan et al., 2021) argue that MI has superior predictive performance compared to impute-then-classify.  Graham (2009) find that MI performs well even with less data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets, and is highly unstable. This can likely be explained by the complexity-stability trade-offs: MI uses a more complex model class and is able to minimize the empirical loss, but due to this very complexity it shows a large variation in predictions under small train set perturbations.
Best performing methods on single- and multi-mechanism missingness are different.
A key novelty in our study is the evaluation under multi-mechanism missingness. We find that MVI methods such as boostclean, which was not competitive for single-mechanism missingness, performs well under mixed or multi-mechanism missingness — which is a more complex missingness pattern and a more realistic scenario likely to be encountered in practice.
Fairness is highly mechanism-specific.
Zhang and Long (2021b) find that different missing mechanisms lead to different prediction fairness when the imputation method is fixed, meaning that an MVI that leads to model fairness under, say, MAR may not lead to fairness under MNAR. We also observe this for all the datasets in our study.
Model stability depends more on the dataset size and MVI technique than on the missingness scenario.
We find that, for large datasets, even simple statistical imputers can preserve stability. In contrast, for small datasets, only a few MVI techniques preserve stability, while deletion, statistical imputation and more complex ML-based approaches all worsen stability.
Sensitivity to missingness rates.
For predictive performance, models are more sensitive to missingness during test than during train. Fairness is highly sensitive to both train and test missingness rates, and, for most methods, the model is most fair when there is the same rate of missingness in both. There is little effect of increasing test missingness on stability, but most imputers show worse stability at higher train missingness rates.
No good methods for MNAR missingness.
MNAR is theoretically the hardest setting to model, and we empirically find that most state-of-the-art MVI methods do not attain predictive performance of the model trained on clean data. Further, our analysis of missingness shift indicates that training under MNAR is highly sensitive to missingness rates, and even methods that are trained on MNAR but evaluated on “simpler” test missingness (MCAR and MAR) perform poorly. Notably, under multi-mechanism missingness, seeing a small amount of MCAR and MAR data during training boosts MVI performance. Designing MVI techniques that are effective under MNAR is an open research direction.
Imputation quality and fairness metrics are often insufficiently predictive of the correctness, fairness and stability of downstream models.
Shadbahr et al. (2023) showed that distributional imputation quality metrics are a more reliable indicator of the model’s predictive performance, than discrepancy metrics under single-mechanism missingness. Our findings validate this result for multi-mechanism missingness as well. Further, we uncover trade-offs between predictive performance, fairness and stability. Finally, we find that imputation fairness metrics are not predictive of model fairness.
It’s complicated!
Prior experimental work (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) concluded that there is no definitive answer to the question of which missing value imputation technique is best, according to predictive performance. Our findings support this, and further demonstrate trade-offs between model fairness and accuracy, and model stability and accuracy. This is not bad news, but rather an acknowledgement of the inherent complexity of the task of learning from incomplete data, and a call for more rigorous and holistic evaluation protocols— of the kind demonstrated in this work — in order to select the most suitable MVI technique for a given task.
8. Conclusions and Future Work
Conclusions. We presented Shades-of-Null, a benchmark for responsible missing value imputation. The key novelty of our work is in evaluating predictive performance of models trained on imputed data alongside fairness and stability. Another novelty is our modeling of realistic missingness scenarios that go beyond Rubin’s classic settings of MCAR, MAR and MNAR to include multi-mechanism missingness and missingness shift. We showcased results of an extensive study, with 20,952 experimental pipelines, that uses our benchmark to evaluate performance of a wide range of missing value imputation methods in realistic scenarios. We observed interesting trends, and identified trade-offs between evaluation metrics.
Limitations.
We started with clean datasets, carefully designed meaningful missingness scenarios, and simulated them using error injectors. A limitation of our work is that we evaluate performance on synthetically generated — rather than on naturally occurring — missingness. This is a common limitation of work in this area: we often do not have a clean ground truth to compare against, and so we make the assumption that injecting synthetic errors will only exacerbate the effect of existing (unknown) errors. With this assumption, any degradation in performance we see is an underestimate of the possible degradation we can expect by comparing with an ideal clean ground truth.
Future work.
While predictive performance, fairness and stability can be seen as orthogonal dimensions of performance, we have uncovered trade-offs among them that arise due to missingness in the data, the choice of a missing value imputation method, and, in some cases, the architecture of the model trained on the data post-imputation. This, coupled with the finding that there is no universally best-performing imputation method, motivates the development of holistic metrics, evaluation methods, and imputation techniques that are grounded in the lifecycle of data collection, processing and use. That is, much exciting work remains in responsible data engineering, of which missing value imputation is an important dimension, and to which our work makes a contribution.
Appendix Overview
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the benchmark implementation, which include a broad description of the missing value imputation techniques used in our study and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair benchmark (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix C.1 presents the imputation quality of all MVI techniques across all datasets from the perspectives of accuracy, distributional alignment, and fairness, while Appendix C.2 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4. Finally, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 6. Specifically, Appendix D.2 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.3 explores further experimental results on model performance under missingness shift with variable train and test error rates.
Appendix A Additional details on benchmark implementation
A.1. Missing Value Imputation Techniques
In our benchmark, we evaluate 8 MVI techniques, classified into 5 main categories based on the taxonomy presented in (Emmanuel et al., 2021). In each category, we selected one or two SOTA methods for our comparison. Each technique is capable of imputing both categorical and numerical features. All configurations and hyperparameters for the MVI techniques described below are specified in the configuration module within the source code.
  1. (1)
    deletion: We simply drop the rows with missingness indicators.
  2. (2)
    Statistical techniques:
    1. (a)
      median-mode: We impute missing values with the median of the complete case for numerical columns and mode (most frequently occurring value) for categorical columns.
    2. (b)
      median-dummy: We impute missing values with the median of the complete case for numerical columns and assign a (new) dummy category for nulls in categorical columns.
  3. (3)
    Machine learning approaches:
    1. (a)
      miss-forest (supervised) (Stekhoven and Bühlmann, 2011): This approach iteratively trains a RF model on a set of clean samples (no missingness) and predicts the missing values.
    2. (b)
      clustering (unsupervised) (Gajawada and Toshniwal, 2012): We assign the missing value to a cluster, based on the distance from the cluster center, and then impute the missing value with the mean or mode of the data points in that cluster. [^7] We used the k-prototypes implementation from https://github.com/nicodv/kmodes We implemented this method ourselves, as no existing solution supports null imputation based on clustering for data with both numerical and categorical features.
  4. (4)
    Deep learning approaches:
    1. (a)
      datawig (Biessmann et al., 2019): This method implements deep learning modules combined with neural architecture search and end-to-end optimization of the imputation pipeline.
    2. (b)
      auto-ml (Jäger et al., 2021): This approach uses the AutoML library autokeras (Jin et al., 2019) to implement the discriminative deep learning imputation method. For categorical columns, autokeras’ StructuredDataClassifier is employed, while StructuredDataRegressor is utilized for numerical columns. These classes manage data encoding and optimize model architecture and hyperparameters autonomously.
  5. (5)
    Multiple imputation (joint) using boostclean (Krishnan et al., 2017): This approach treats the error correction task as a statistical boosting problem where a set of weak learners are composed into a strong learner. To generate the weak learners, BoostClean iteratively selects a single imputation technique, applies it to a training set (with missingness), and fits a new model on the newly imputed training set.
A.2. Corrections and Enhancements in Benchmarks and MVI Techniques
Prior to constructing our benchmark, we conducted a thorough analysis of the strengths and weaknesses of existing state-of-the-art benchmarks and MVI techniques by reviewing their codebases. Unfortunately, we identified several methodological flaws and code bugs in some of them. Consequently, alongside creating our software novelty in the benchmark, we dedicated considerable effort to correcting these bugs in the null imputers and enhancing existing benchmarking approaches to ensure the development of a truly fair benchmark.
Although our benchmark pursued different goals, (Abdelaal et al., 2023) served as the primary benchmark against which we compared our work. The authors made significant contributions, establishing a benchmark that set a high standard in the field. However, we identified several critical concerns in their codebase that we aimed to address. Firstly, their comparison did not include hyper-parameter tuning for null imputers, and they did not utilize seeds to control randomization in the null imputers. In contrast, we meticulously tuned hyper-parameters for each null imputer in our study, implemented controlled randomization throughout the entire pipeline using seeds, and rigorously tested the benchmark with both unit and integration tests to ensure reproducibility. Secondly, a significant methodological flaw in their benchmark was that null imputers were fitted on the entire dataset without initially splitting it into training and test sets. The dataset split was introduced only during the model training stage, resulting in a clear leakage when null imputers were fitted on samples included in the test set. This issue likely arose because imputers like miss-forest and datawig lack built-in interfaces for fitting only on a training set and then transforming both training and test sets. Consequently, the authors chose to use these imputers without modification. In contrast, our benchmark aimed to ensure a fair comparison of MVI techniques as they would be used in production systems. To achieve this, we fixed and extended existing MVI techniques, introducing another software innovation in our benchmark.
In our study, we standardized all 8 MVI techniques discussed in Section 3.2 to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameters tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable). Below are the enhancements we implemented for MVI techniques used in our benchmark:
  • clustering: We implemented this method ourselves to meet our imputation requirements, as existing solutions do not support null imputation based on clustering for datasets containing both numerical and categorical features. For this, we used the k-prototypes as a distance function in the imputer. [^8] https://github.com/nicodv/kmodes
  • miss-forest (Stekhoven and Bühlmann, 2011): The main concern with the original implementation was that the null imputer was fitted within a transform method. This meant that, when applying the transform method to a test set, the null imputer would be fitted on testing samples as well. To adhere to our imputation requirements, we moved the predictor fitting to a fit method and reused the fitted predictors in the transform method. Additionally, we introduced hyper-parameter tuning for the base classifier and regressor, which was lacking in the original implementation.
  • datawig (Biessmann et al., 2019): To align this technique with our imputation criteria, we modified the complete method from the
    SimpleImputer, which imputes null values across all categorical and numerical columns in the dataset. We adjusted the methodological approach in the source code and reused built-in fit and transform methods from the SimpleImputer to achieve this.
  • auto-ml: We adopted a conceptual idea from (Jäger et al., 2021) and used their code as a foundation for our implementation. A significant limitation in their source code was its capability to handle only one column with null values at a time. To enhance its functionality, we expanded the method to support imputation across multiple categorical and numerical columns. Additionally, we incorporated ideas inspired by miss-forest (Stekhoven and Bühlmann, 2011), specifically leveraging initial statistical estimates for each column containing null values to train classifiers and regressors used for imputation.
  • boostclean (Krishnan et al., 2017): We enhanced this method by introducing the capability to tune the internal hyper-parameters of boostclean used for statistical boosting. Furthermore, we implemented the ability to fine-tune the hyper-parameters of the underlying models trained on each dataset that was imputed using MVI techniques listed in boostclean.
In summary, our benchmark offers users not only a robust evaluation framework but also incorporates the refined MVI techniques developed as part of our work.
Appendix B Additional Experimental Details for Reproducibility
B.1. Computing Infrastructure
Our large-scale experimental study incorporated 8 null imputation techniques, 5 model types, 6 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline. This configuration led to an overwhelming total of 20,952 experimental pipelines. The complexity was further compounded by the necessity to tune both null imputers and ML models within each pipeline, employing a bootstrap of 50 estimators to assess model stability. Depending on the dataset size, null imputation technique, and model type, the execution time for a single experimental pipeline ranged from 3 minutes to 16.5 hours. Executing such a large number of pipelines would be impossible without a suitable experimental environment comprising a high-performance computing (HPC) cluster for execution and a 3-node MongoDB cluster for storing experimental results.
Our benchmark controller provides a useful separation of experimental pipelines, and the MongoDB cluster gives the ability to store results independently from a local file system. To manage the extensive computational demands, we utilized several optimizations (detailed in Section 3.4) and the SLURM software system, which allowed us to execute up to 100 simultaneous jobs on the HPC cluster. To run experimental pipelines, each SLURM job was allocated 12-24 physical cores or one GPU card, along with 16-96 GB of RAM, depending on the dataset size, imputation method, and model type. For null imputation using datawig, we employed an RTX8000 NVIDIA GPU card with 48 GB of GPU memory, while all other tasks were handled by a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz processor.
To effectively store the results of the experiments in the MongoDB cluster, we designed a data model with unique GUIDs for each experimental pipeline. This model comprehensively encompasses all facets of each experiment: dataset names, all used seeds, imputer and model hyper-parameters, and more. Since we have multiple options of input variables for each stage of the experimental pipeline, we needed to consolidate and unify the data to enable effective analysis of the results of the experiments. Hence, we established GUIDs for each layer of the experimental pipeline, encompassing identifiers for each imputation stage, model tuning stage, model profiling stage, and experimental session. This data model helps us access data on different levels, unify the structure of all experiments, increase the safety of aggregations and result interpretation, and simplify the creation of visualizations.
B.2. Extended Methodology for Simulating Missingness
Similarly to Section 3.1, this subsection outlines realistic missingness scenarios based on intrinsic dataset properties. Tables 4-9 provide the demographic composition of all six datasets, expressed through the proportions and base rates of each protected group. Figures 13-17 display the Spearman correlation coefficients between covariates and the target variable, as well as feature importance derived using scikit-learn’s built-in functionality for datasets not covered in Section 3.1. For datasets with numerous features, plots include only the most correlated features. These proportions, base rates, correlation coefficients, and feature importance guided the conditions of our missingness scenarios, aiming to reflect realistic instances of non-response caused by disparate access, distrust, or procedural injustice, where disadvantaged groups exhibit more missing values than privileged groups. Tables 10-14 detail the missingness scenarios for the folk, bank, heart, law-school, and german datasets, respectively.
Table 4. Table 4: Proportions and Base Rates for german.
overall sex_priv sex_dis age_priv age_dis sex&age_priv sex&age_dis
Proportions 1.0 0.69 0.31 0.81 0.19 0.895 0.105
Base Rates 0.7 0.723 0.648 0.728 0.579 0.717 0.552
Table 5. Table 5: Proportions and Base Rates for folk.
overall sex_priv sex_dis race_priv race_dis sex&rac1p_priv sex&race_dis
Proportions 1.0 0.511 0.489 0.678 0.322 0.829 0.171
Base Rates 0.35 0.422 0.274 0.389 0.267 0.374 0.232
Table 6. Table 6: Proportions and Base Rates for law-school.
overall male_priv male_dis race_priv race_dis male&race_priv male&race_dis
Proportions 1.0 0.561 0.439 0.841 0.159 0.917 0.083
Base Rates 0.89 0.899 0.878 0.921 0.723 0.906 0.713
Table 7. Table 7: Proportions and Base Rates for diabetes.
overall gender_priv gender_dis
Proportions 1.0 0.621 0.379
Base Rates 0.291 0.272 0.321
Table 8. Table 8: Proportions and Base Rates for bank.
overall age_priv age_dis
Proportions 1.0 0.955 0.045
Base Rates 0.117 0.106 0.341
Table 9. Table 9: Proportions and Base Rates for heart.
overall gender_priv gender_dis
Proportions 1.0 0.35 0.65
Base Rates 0.5 0.505 0.497
Table 10. Table 10: Missingness scenarios for an error rate of 30% for folk. WKHP, AGEP are numerical columns; SCHL, MAR are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR WKHP, AGEP, SCHL, MAR N/A 0.3 0.3
MAR WKHP, SCHL SEX 0.2 (female) 0.1 (male)
MAR, AGEP RAC1P 0.2 (non-white) 0.1 (white)
MNAR MAR MAR 0.25 (not married) 0.05 (married)
WKHP WKHP 0.25 (¡ 40) 0.05 ($\geq 40$)
AGEP AGEP 0.25 (¿ 50) 0.05 ($\leq 50$)
SCHL SCHL 0.25 (¡ 21) 0.05 ($\geq 21$)
(a) Correlation with label
(b) Feature importance
Figure 13. EDA for designing missingness scenarios in folk.
Table 11. Table 11: Missingness scenarios for an error rate of 30% for bank. balance, campaign are numerical columns; education, job are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR balance, campaign, education, job N/A 0.3 0.3
MAR education, job age 0.18 ($\geq 30$) 0.12 (¡ 30)
balance, campaign marital 0.2 (single) 0.1 (married)
MNAR education education 0.2 (tertiary) 0.1 (secondary)
[…]
balance balance 0.2 ($\leq 1000$) 0.1 (¿ 1000)
campaign campaign 0.2 ($\leq 1$) 0.1 (¿ 1)
(a) Correlation with label
(b) Feature importance
Figure 14. EDA for designing missingness scenarios in bank.
Table 12. Table 12: Missingness scenarios for an error rate of 30% for heart. weight, height are numerical columns; cholesterol, gluc are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR weight, height, cholesterol, gluc N/A 0.3 0.3
MAR weight, height gender 0.2 (female) 0.1 (male)
cholesterol, gluc age 0.2 ($\geq 50$) 0.1 (¡ 50)
MNAR weight weight 0.25 ($\geq 75$) 0.05 (¡ 75)
height height 0.2 (¡ 160) 0.1 ($\geq 160$)
cholesterol cholesterol 0.16 (not normal) 0.14 (normal)
gluc gluc 0.12 (not normal) 0.18 (normal)
(a) Correlation with label
(b) Feature importance
Figure 15. EDA for designing missingness scenarios in heart.
Table 13. Table 13: Missingness scenarios for an error rate of 30% for law-school. zfygpa, ugpa are numerical columns; fam_inc, tier are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR zfygpa, ugpa, fam_inc, tier N/A 0.3 0.3
MAR ugpa, zfygpa male 0.2 (0) 0.1 (1)
am_inc, tier race 0.15 (non-white) 0.15 (white)
MNAR ugpa ugpa 0.2 (¡ 3.0) 0.1 ($\geq 3.0$)
zfygpa zfygpa 0.2 ($\leq 0$) 0.1 (¿ 0)
fam_inc fam_inc 0.2 (¡ 4) 0.1 ($\geq 4$)
tier tier 0.2 (¡ 4) 0.1 ($\geq 4$)
(a) Correlation with label
(b) Feature importance
Figure 16. EDA for designing missingness scenarios in law-school.
Table 14. Table 14: Missingness scenarios for an error rate of 30% for german. duration, credit-amount are numerical columns; checking-account, savings-account, employment-since are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR duration, credit-amount, checking-account, savings-account, employment-since N/A 0.3 0.3
MAR savings-account, checking-account, credit-amount age 0.18 ($\leq 25$) 0.12 (¿ 25)
employment-since, duration sex 0.2 (female) 0.1 (male)
MNAR checking-account checking-account 0.25 (no account) 0.05 (not no account)
duration duration 0.25 ($\leq 20$) 0.05 (¿ 20)
savings-account savings-account 0.2 (not no savings account) 0.1 (no savings account)
[…]
credit-amount credit-amount 0.25 (¿ 5000) 0.05 ($\leq 5000$)
(a) Correlation with label
(b) Feature importance
Figure 17. EDA for designing missingness scenarios in german.
Appendix C Additional Experimental Results for Single-mechanism Missingness
C.1. Imputation Quality and Fairness
To better understand the impact of different MVI techniques on model performance, we measure the imputation quality of each imputer before training ML models on the imputed training set. The quality of imputation is evaluated across three dimensions: correctness (using F1 for categorical features and RMSE for numerical features), distributional alignment (using KL divergence of imputed values), and fairness (measuring differences in F1, RMSE, and KL divergence between disadvantaged and privileged groups). The higher values of F1 are the better, the lower values of RMSE and KL divergence are the better, and values of imputation fairness closer to zero are the better.
To reduce the number of plots, we selected one categorical and one numerical features for each dataset based on their Spearman correlation with the target. However, our analysis takes into account the imputation quality measured on all features with nulls. Importantly, these figures do not include boostclean since this is a joint cleaning and model training method that does not provide imputed values, which we could use to measure imputation quality. The x-axis labels include the names of the corresponding features from the dataset for which the imputation quality metric was measured.
Correctness of imputation. Figures 18 and 19 demonstrate the correctness of imputation of all MVI techniques, except boostclean, for all 6 datasets. From Figure 18, we can notice that one of the ML and DL-based imputers (miss-forest, auto-ml, datawig) is the best in terms of F1 for categorical columns under MCAR and MAR missingness mechanisms. Additionally, Figure 19 shows that all the three ML and DL-based imputers are dominant in terms of RMSE for numerical columns compared to statistical MVI techniques. However, these significant differences in F1 and RMSE do not translate into substantial differences in downstream model performance, as measured by F1 in Section 4.1. This observation further supports our statement made in Section 5.4 that ”imputation quality metrics are not very meaningful because a bad data imputation method can be overcome by a strong learner, and so bad performance on imputation quality does not necessarily mean than the MVI will lead to poor predictive performance.”
Distributional alignment of imputation. Figures 20 and 22 illustrate the distributional alignment of imputation of all MVI techniques except boostclean, for all 6 datasets. The plots clearly demonstrate that ML and DL-based imputers, except clustering, outperform statistical approaches in terms of distributional alignment across all missingness mechanisms. In Section 5.4, we highlight scenarios where distributional metrics serve as more reliable indicators than discrepancy metrics. This is because the distributional alignment of imputed features can significantly influence model training outcomes, emphasizing that assessing only the correctness of imputation is often insufficient.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 18. F1 of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 19. RMSE of different imputation strategies (colors in the legend) for the most correlated numerical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 20. KL divergence of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 21. F1 Difference of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 22. KL divergence of different imputation strategies (colors in the legend) for the most correlated numerical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 23. RMSE Difference of different imputation strategies for the most correlated numerical column.
Fairness of imputation. Figures 21, 23, 24, and 25 show the fairness of imputation of all MVI techniques, except boostclean, for all 6 datasets. We report fairness metrics based on an intersectional group (i.e., sex&race) for datasets with two sensitive attributes, and a singular group (i.e., age) for datasets with one sensitive attribute. From Figures 21 and 23, we can notice that statistical MVI techniques are prone to worsen imputation fairness rather than improve it under all missingness mechanisms. Further, comparing these imputation fairness differences with the model fairness differences discussed in Section 4, we can state that good imputation fairness does not guarantee good model fairness.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 24. KL Divergence Difference of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 25. KL Divergence Difference of different imputation strategies (colors in the legend) for the most correlated numerical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
C.2. Additional Accuracy and Fairness Metrics of Predictive Model
In this section, we extend the results discussed in Section 4 by presenting plots for additional popular fairness metrics of downstream model performance. Figures 26, 27, and 28 illustrate Disparate Impact, True Negative Rate Difference, and Selection Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms.
Overall, the plots support the assertion made in Section 4.2 that the impact of MVI on fairness (error-disparity) is highly dataset-specific and is most strongly influenced by the baseline disparity of the model trained on clean (complete, no missingness) data. This implies that it is impossible to determine in advance whether data cleaning will worsen fairness.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 26. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a clean baseline.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 27. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a clean baseline.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 28. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a clean baseline.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 29. Accuracy of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are ordered in increasing order by size. The blue line shows performance of a model trained on clean data.
Appendix D Additional Experimental Results for Missingness Shift
D.1. Stability of Predictive Model
This section supplements the discussion of Section 6.3. It includes the label stability plots that were omitted in the main body of the paper due to space constraints.
Figure 30. Label Stability of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Figure 31. Label Stability of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
D.2. Fixed Train and Test Missingness Fractions
F1 of predictive model. The F1 scores of predictive models trained with various MVI techniques under different train and test missingness conditions, including missingness shift (scenarios S1-9 from Table 1), are presented in Figure 34. The plot demonstrates that the F1 scores of predictive models are sensitive to missingness shifts across different datasets, as also discussed in Section 6. The impact of missingness shift is most significant for MCAR train & MNAR test and MAR train & MNAR test scenarios on the diabetes, german, and heart datasets. This observation aligns with the statement in (Shadbahr et al., 2023) that MNAR is more complex to model than MCAR and MAR. Additionally, miss-forest emerges as the most robust technique in handling missingness shifts across all datasets and settings.
Figure 32 further supports these findings by showing the performance of different model types in terms of F1 scores on diabetes after applying various ML and DL-based null imputers.
(a) MissForest
(b) AutoML
(c) K-Means Clustering
(d) Deletion
Figure 32. F1 of different models (colors in legend) on the diabetes dataset, for best and worst performing MVI techniques (subplots).
Stability of predictive model. The stability of various top-performing models using different MVI techniques under missingness shift is illustrated in Figure 35. The plot indicates that missingness shift does not significantly impact model stability for most MVI techniques, compared to the impact observed in F1 scores. However, the plot shows that the choice of MVI technique can lead to notable differences in model stability for small-sized datasets, whereas changes in stability are less significant for datasets with more than 15K rows. This suggests that ML models exhibit better stability with larger datasets. Additionally, the plot supports our observation in Section 4.3 that the state-of-the-art multiple imputation technique boostclean shows lower model stability across different datasets compared to other MVI techniques, even if it holds good model accuracy.
Figure 33 further demonstrates the effect of model type on stability for the diabetes dataset.
(a) MissForest
(b) AutoML
(c) K-Means Clustering
(d) Deletion
Figure 33. Stability of different models (colors in legend) on the diabetes dataset, for best and worst performing MVI techniques (subplots).
(a) Diabetes (905 samples)
(b) German (1000 samples)
(c) Folk (15,000 samples)
(d) Law School (20,798 samples)
(e) Bank (40,004 samples)
(f) Heart (70,000 samples)
Figure 34. F1 of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-9)
(a) Diabetes (905 samples)
(b) German (1000 samples)
(c) Folk (15,000 samples)
(d) Law School (20,798 samples)
(e) Bank (40,004 samples)
(f) Heart (70,000 samples)
Figure 35. Label stability of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-9)
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 36. The plot reveals that the impact of missingness shifts on fairness varies depending on the missingness type of the test set. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 37-54. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across all ML and DL-based imputers and all six datasets.
(a) Diabetes and Random Forest
(b) Law School and Logistic Regression
(c) Bank and Gradient Boosted Trees
Figure 36. Fairness, measured by True Positive Rate Difference, of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-9)
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 37. Disparate Impact of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 38. True Negative Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 39. Selection Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 40. Disparate Impact of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 41. True Negative Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 42. Selection Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 43. Disparate Impact of different models (colors in legend) on the folk dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 44. True Negative Rate Difference of different models (colors in legend) on the folk dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 45. Selection Rate Difference of different models (colors in legend) on the folk dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 46. Disparate Impact of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 47. True Negative Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 48. Selection Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 49. Disparate Impact of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 50. True Negative Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 51. Selection Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 52. Disparate Impact of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 53. True Negative Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots).
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 54. Selection Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots).
D.3. Variable Train and Test Missingness Fractions
Fairness under train set missingness. Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
The plots for additional fairness metrics support the assertion in Section 6.2 that null imputation methods are sensitive to the train missingness rate in terms of fairness. However, the impact varies across different fairness metrics. The highest deviation from the baseline value of each fairness metric is observed in the subfigures for MAR train sets at a 30% error rate, matching the test error rate. For instance, the effect on TNRD is similar to that on TPRD in Section 6.2, with most imputers showing an improvement in TNRD, whereas Disparate Impact and Selection Rate Difference worsen as the train error rate increases.
Fairness under test set missingness. Figures 58, 59, and 60 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.
The plots reinforce the assertion in Section 6.2 that the trend depends significantly on the test missingness type. For example, the subfigures for MCAR train & MNAR test and MAR train & MNAR test show the largest deviations with increasing test error rates for all three fairness metrics, while plots for other settings remain relatively flat. Interestingly, when null imputers and models are fitted on the MNAR train set, model performance on the MNAR test set across different test error rates do not change significantly compared to MCAR train & MNAR test and MAR train & MNAR test settings. This indicates that the missingness mechanism in the train set can significantly affect model fairness on the MNAR test set. Additionally, miss-forest and auto-ml proved to be the most robust MVI techniques against test set missingness shifts under all conditions.
Figure 55. True Negative Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
Figure 56. Disparate Impact of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
Figure 57. Selection Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
Figure 58. True Negative Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
Figure 59. Disparate Impact of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.
Figure 60. Selection Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of a clean baseline.